OpenStudy (anonymous):

find the integral:

OpenStudy (anonymous):

\[\int\limits_{-3}^{3} \sqrt{-1/9 \times X^2 + 1} dx\]

OpenStudy (chaise):

Do you just want an answer or how to work it out?

OpenStudy (anonymous):

my bad, no square root

OpenStudy (anonymous):

and the answer is supposed to be 16 but I keep getting 4...

OpenStudy (chaise):

\[\int\limits_{-3}^{3} \frac{-1}{9}x^2+1\] Is that what you're asking?

OpenStudy (anonymous):

yup

OpenStudy (chaise):

My calculator says 4, perhaps the answer you are provided with is incorrect :)

OpenStudy (anonymous):

or maybe im doing something wrong, the original question is to find the area of a solid whose base is enclosed by 1/9 x^2 + y^2=1 with square cross sections

OpenStudy (chaise):

What you have provided is the equation for an ellipse. You are trying to find the area of the base of the ellipse, which is given by that equation. You can find the area of this ellipse with a formula. I hope this is what the question is asking.

OpenStudy (anonymous):

no it's a solid, where the base is an ellipse and the cross section is square

OpenStudy (chaise):

It's a little counter intuitive, but what you have to imagine is this: http://imageshack.us/photo/my-images/337/hellooooooooooooooooooo.jpg/